# Cascaded Multi-path Shortcut Diffusion Model for Medical Image Translation

### v1

@@ Abstract
Image Translation, Diffusion Model, Uncertainty, Cascade Framework
@@ 1 Introduction
Figure 1: Illustration of previous I2I diffusion model generation process. Starting the reverse process with different noise initialization leads to divergent translation results.
@@
Figure 2: The overall workflow of our proposed Cascade Multi-path Shortcut Diffusion Model (CMDM). CMDM consists of a one-step inference model (green) and cascades of MPD block (grey). Each MPD block consists of multiple shortcut reverse paths starting with a prior image with different noise. The cascades are connected with residual averaging operations.
@@
On the other hand, I2I diffusion models have been recently developed and show superior performance than GANs. For general-purpose I2I with DM, [25] proposed a unified framework, Palette, which adds conditional image inputs to the previously developed Denoising Diffusion Probabilistic Model (DDPM [10]), thus enabling the I2I functionality of DDPM. To reduce the randomized initialization process and improve the stability in I2I DM, direct bridging diffusion methods have been investigated. Notably, [15] developed a Brownian Bridge Diffusion Model (BBDM) that learns the translation between two domains directly through the bidirectional diffusion process, i.e. Brownian Bridge, rather than a conditional generation process. Similarly, [16] proposed a Schrodinger Bridge I2I Diffusion Model (I2SB) that directly learns the nonlinear diffusion processes between two domains. Both had shown improved I2I performance in natural image translation tasks. Similar to I2I GANs, these DM methods have been applied in medical imaging. For example, [20] utilized DDPM to synthesize artificial histopathology images with rare cancer subtypes to mitigate data imbalance problems for medical data. [18] proposed to translate CT into MRI with conditional DDPM and score-matching models. The forward and backward diffusion processes are guided by T2 MRI. [8] proposed to perform brain PET image denoising with MRI as prior information to improve image quality. [7] utilized a contextual contained network in the DM to improve the LDCT denoising. Furthermore, 2D DMs have also been employed for 3D translation tasks, including low-count PET image denoising [35], CT reconstruction [4], and MRI super-resolution and reconstruction [14]. Direct extension to 3D DM were also explored [23]. However, there are several unique challenges of DM for I2I. First, those methods require iterating over a large number of steps in the reverse process, and most methods start the generation with pure random noise [25, 18, 8, 35, 4, 14]. This protocol not only significantly slows down the translation speed, but could also lead to diverged and sub-optimal translation results if different random noise initialization were used in the input when running multiple reverse runs (Figure 1). Even though direct bridging methods [15, 16] are translation deterministic given that no random noise input is used, they still require a large number of reverse iteration steps. Another challenge of the deterministic translation is that they also cannot generate translation uncertainty maps which is crucial for medical images, since the model’s prediction error can be used to pinpoint problem areas or give clinicians more information [28, 13, 33]. It is then a unique advantage of the stochastic sampling process of the conditional DDPM [25] to obtain the uncertainty map through running the DM repeatedly with multiple random noises [33]. Therefore, it is desirable to develop an I2I DM method that can generate high-quality converged translation results with a reduced number of required iterations, while also being able to provide translation uncertainty estimation.
@@ 2.1 Cascaded Multi-path Shortcut Diffusion Model
Training: Let us denote the input image as $x$ and the translation target as $y_{0}$. For the prior image generation part, we utilized a generative network, i.e. UNet [24], that aims to predict $y_{0}$ from $x$. The network can be trained in an adversarial fashion using both a pixel-wise L2 loss
@@
Figure 3: Qualitative comparison of translation results and corresponding error map from different methods. Examples from DE X-ray soft-tissue generation (Left), Sparse-view CT reconstruction (Middle), and MRI T1-to-T2 synthesis are shown. The image quality metrics of each sample are indicated at the bottom left of the images.
Inference: Once the prior image generation network $f_{prior}(\cdot)$ and the conditional diffusion network $f_{dm}(\cdot)$ are converged from training, we can use them in CMDM for image translation. The overall inference pipeline of CMDM is illustrated in Figure 2. Instead of starting the reverse process from a standard normal distribution $\mathcal{N}(y_{T}|0,I)$ at $T$, the reverse process starts at a pre-defined time point $t_{s}\in[0,T]$ with
@@
UNet $39.76\pm 2.36$$984\pm.003$$606\pm.071$ $41.33\pm 3.18$$988\pm.003$$571\pm.066$ 0.013
cGAN $39.82\pm 2.37$$985\pm.003$$603\pm.072$ $41.36\pm 3.17$$988\pm.003$$572\pm.065$ 0.013
Palette v1 $42.89\pm 2.34$$987\pm.002$$390\pm.047$ $43.06\pm 3.16$$989\pm.002$$373\pm.042$ 13.670
Palette v2 $43.11\pm 2.34$$988\pm.002$$382\pm.045$ $43.47\pm 3.13$$990\pm.002$$363\pm.043$ 273.420
I2SB $43.18\pm 2.35$$988\pm.002$$381\pm.045$ $43.49\pm 3.14$$990\pm.002$$367\pm.043$ 14.551
BBDM $43.08\pm 2.35$$988\pm.002$$382\pm.044$ $43.52\pm 3.13$$989\pm.002$$359\pm.043$ 15.121
Ours $\mathbf{44.27\pm 2.33}$$\mathbf{.991\pm.002}$$\mathbf{.369\pm.041}$ $\mathbf{44.58\pm 3.16}$$\mathbf{.992\pm.003}$$\mathbf{.348\pm.038}$ 154.663
@@
UNet $44.11\pm 1.38$$977\pm.004$$372\pm.047$ $46.32\pm 1.27$$981\pm.004$$315\pm.040$ 0.006
cGAN $44.13\pm 1.39$$978\pm.004$$370\pm.047$ $46.35\pm 1.28$$981\pm.004$$314\pm.040$ 0.006
Palette v1 $44.96\pm 1.24$$980\pm.003$$321\pm.041$ $46.75\pm 1.26$$987\pm.004$$310\pm.039$ 8.863
Palette v2 $45.56\pm 1.24$$981\pm.003$$318\pm.040$ $46.95\pm 1.25$$988\pm.003$$308\pm.038$ 177.202
I2SB $45.86\pm 1.26$$982\pm.003$$317\pm.039$ $46.91\pm 1.26$$989\pm.003$$309\pm.039$ 9.561
BBDM $45.73\pm 1.24$$981\pm.003$$318\pm.040$ $46.96\pm 1.26$$989\pm.003$$309\pm.038$ 9.987
Ours $\mathbf{46.42\pm 1.22}$$\mathbf{.986\pm.003}$$\mathbf{.302\pm.039}$ $\mathbf{47.02\pm 1.25}$$\mathbf{.990\pm.003}$$\mathbf{.299\pm.038}$ 108.821
@@
UNet $27.17\pm 1.56$$885\pm.042$$2$051 $27.38\pm 1.59$$891\pm.046$$216\pm.052$ 0.006
cGAN $27.19\pm 1.58$$887\pm.044$$220\pm.052$ $27.41\pm 1.58$$891\pm.047$$217\pm.053$ 0.006
Palette v1 $27.52\pm 1.57$$890\pm.044$$218\pm.051$ $27.68\pm 1.54$$897\pm.046$$210\pm.052$ 8.863
Palette v2 $27.68\pm 1.55$$891\pm.043$$211\pm.051$ $27.79\pm 1.52$$899\pm.044$$206\pm.051$ 177.202
I2SB $27.85\pm 1.56$$892\pm.043$$208\pm.051$ $27.89\pm 1.54$$898\pm.043$$208\pm.052$ 9.561
BBDM $27.88\pm 1.56$$892\pm.043$$207\pm.051$ $27.86\pm 1.53$$899\pm.045$$207\pm.051$ 9.987
Ours $\mathbf{27.93\pm 1.54}$$\mathbf{.898\pm.042}$$\mathbf{.202\pm.051}$ $\mathbf{27.98\pm 1.54}$$\mathbf{.901\pm.044}$$\mathbf{.201\pm.051}$ 108.821
@@ 2.2 Dataset Preparation
Figure 4: Ablative studies on the reverse starting time (Left), the number of paths (Middle), and the number of cascades (Right). DE X-ray soft-tissue image generation and 1/6 SVCT reconstruction were utilized for these studies. Peak performances were annotated on the plots with the corresponding image quality metric, i.e. PSNR.
@@ 3 Experimental Results
Figure 3 shows qualitative comparisons between previous state-of-the-art and our methods. Examples from the DE X-ray dataset, SVCT reconstruction dataset, and MRI translation dataset are illustrated. For the DE X-ray example (left two columns), we can see all the previous translation methods can generate reasonable soft-tissue images, i.e. rib-suppression images, from the standard X-ray image. While generating visually plausible results with a PSNR of $44.74$dB from cGAN, it still suffers from relatively inaccurate quantification as indicated by the error map. On the other hand, we can see the previous diffusion-based methods, e.g. Palette and BBDM, both achieved significantly better translation as compared to cGAN with PSNR improving to $46.05$dB, with much fewer pixel-wise errors indicated by the error maps. On the last row, we can find that our CMDM further improved over the previous diffusion-based methods with PSNR reaching over $47$dB, where further reduced pixel-wise error can be found in the cardiac and lung regions. Similarly, for the SVCT example (middle two columns), cGAN can reasonably suppress the streak artifact in the input FBP SVCT image. However, significant residual errors can be found in the femoral head and pelvic bone regions. On the other hand, we can see that the previous diffusion-based methods can suppress these errors, with PSNR reaching close to $46$dB. With CMDM combining cGAN and Diffusion, we can see that our translation result further reduces the overall error and improves the image quality to PSNR=$46.23$. Similar observations can be found for the T1-to-T2 translation example in the last two columns.
@@
The quantitative comparisons were summarized in Table 1. Similar to the observations from the visualizations, we can see the traditional CNN-based approaches generally under-performed the diffusion-based approaches. For example, the cGAN only achieved an average PSNR=$39.82$ and MAE=$0.603$ for the soft-tissue image translation, while the single reverse path Palette, i.e. Palette v1, significantly outperformed it with PSNR=$42.89$ and MAE=$0.390$. Running multiple reverse paths of Palette and averaging the outputs, i.e. Palette v2, leads to improved performance which reached similar performances of I2SB and BBDM with PSNR=$43.11$ and MAE=$0.382$. In the last row, our CMDM achieved an average PSNR=$44.27$ and MAE=$0.369$ that significantly outperformed all the previous baseline methods. Comparing the soft-tissue image translation task to the bone image translation task, the CMDM had slightly higher performance on the latter task since the bone image without complex soft-tissue texture can be relatively easier to generate as compared to the soft-tissue image. For the inference speed in DE X-ray applications, I2SB and BBDM with a single reverse path took an average of $14.55$ and $15.12$ seconds, respectively. CMDM with the best performance took an average of $154.66$ seconds per inference since multiple shortcut reverse paths are needed. Similar to the quantitative results for the DE X-ray, we found our CMDM consistently outperformed previous CNN and diffusion-based baseline methods for both the SVCT reconstruction applications and the MRI inter-modality translation applications.
@@
Figure 5: Examples of CMDM’s uncertainty estimation for DE X-ray soft-tissue image generation (Left) and MRI T1-to-T2 synthesis (Right). The relationship plots between the absolute error (bottom left) and the uncertainty (bottom right) were shown as well. Positive correlations with $R>0.75$ were found for both cases.
@@
We conducted ablative studies for the hyper-parameters in CMDM, including the reverse starting time, the number of shortcut paths, and the number of cascades. The results for the DE X-ray and SVCT are summarized in Figure 4. First, for the reverse starting time, we can see that setting $t_{s}$ to around $200$ yields the best performance, and the performance starts to degrade if we further increase it. It is worth noticing that using $t_{s}=200$ here not only yields the best performance but allows us to reduce the inference time by about 5 times as compared to the previous diffusion methods that start at $t=1000$ or beyond. Second, for the number of shortcut paths, we can see that the performance increases as we use an increasing number of paths. The performance started to converge when $20$ paths were used. Because the inference time increases linearly as we increase the number of paths, we chose the converging point $N_{p}=20$ in our method. Finally, for the number of cascades, we found that the performance gradually boosted as the number of cascades increased. However, peak performance was reached when $N_{c}=3$, and the inference started to overfit, leading to degraded translation performance.

### v2

@@ Abstract
MSC:
[…]
journal: Medical Image Analysis
@@ 1 Introduction
Fig. 1: Illustration of previous I2I diffusion model generation process. Starting the reverse process with different noise initialization leads to divergent translation results.
@@
Fig. 2: The overall workflow of our proposed Cascade Multi-path Shortcut Diffusion Model (CMDM). CMDM consists of a one-step inference model (green) and cascades of MPD block (grey). Each MPD block consists of multiple shortcut reverse paths starting with a prior image with different noise. The cascades are connected with residual averaging operations.
On the other hand, I2I diffusion models have been recently developed and show superior performance than GANs. For general-purpose I2I with DM, [33] proposed a unified framework, Palette, which adds conditional image inputs to the previously developed Denoising Diffusion Probabilistic Model (DDPM [13]), thus enabling the I2I functionality of DDPM. To reduce the randomized initialization process and improve the stability in I2I DM, direct bridging diffusion methods have been investigated. Notably, [21] developed a Brownian Bridge Diffusion Model (BBDM) that learns the translation between two domains directly through the bidirectional diffusion process, i.e. Brownian Bridge, rather than a conditional generation process. Similarly, [23] proposed a Schrodinger Bridge I2I Diffusion Model (I2SB) that directly learns the nonlinear diffusion processes between two domains. Both had shown improved I2I performance in natural image translation tasks. Similar to I2I GANs, these DM methods have been applied in medical imaging.
@@
For example, [27] utilized DDPM to synthesize artificial histopathology images with rare cancer subtypes to mitigate data imbalance problems for medical data. [25] proposed to translate CT into MRI with conditional DDPM and score-matching models. The forward and backward diffusion processes are guided by T2 MRI. [10] proposed to perform brain PET image denoising with MRI as prior information to improve image quality. [9] utilized a contextual contained network in the DM to improve the LDCT denoising. Furthermore, 2D DMs have also been employed for 3D translation tasks, including low-count PET image denoising [45], CT reconstruction [5], and MRI super-resolution and reconstruction [20]. Direct extension to 3D DM were also explored [31]. However, there are several unique challenges of DM for I2I. First, those methods require iterating over a large number of steps in the reverse process, and most methods start the generation with pure random noise [33, 25, 10, 45, 5, 20]. This protocol not only significantly slows down the translation speed, but could also lead to diverged and sub-optimal translation results if different random noise initialization were used in the input when running multiple reverse runs (Figure 1). Even though direct bridging methods [21, 23] are translation deterministic given that no random noise input is used, they still require a large number of reverse iteration steps. Another challenge of the deterministic translation is that they also cannot generate translation uncertainty maps which is crucial for medical images, since the model’s prediction error can be used to pinpoint problem areas or give clinicians more information [36, 16, 42]. It is then a unique advantage of the stochastic sampling process of the conditional DDPM [33] to obtain the uncertainty map through running the DM repeatedly with multiple random noises [42]. Therefore, it is desirable to develop an I2I DM method that can generate high-quality converged translation results with a reduced number of required iterations, while also being able to provide translation uncertainty estimation.
@@
In the field of medical field, a wide spectrum of imaging techniques provides accurate visualization of anatomical structures of human bodies, including ultra sound, computed tomography (CT), magnetic resonance imaging (MRI), and positron emission tomography (PET). Each of them has unique advantages, and can be used for different purposes. For example, T1 MRI highlights soft tissues and is extensively applied to brain imaging. PET shows metabolic function of human and can help clinicians to locate tumors. However, post-processing is an essential step to extract accurate information from these imaging modalities for diagnosis. For example, the intrinsic imprecision of various image acquisition methods; therefore, noise reduction is necessary in improving image quality. The hybrid information from different imaging modalities often help the diagnosis; thus, translating images to another modality and generating the paired images for multiple modalities has been extensively studied [1]. For example, MR images can be translated into corresponding CT images for the purpose of PET attenuation correction. During image acquisition, sparse-measurement conditions (sparse view or limited angles) might cause challenges for accurate reconstruction, so the complementary step is usually needed before using the conventional reconstruction for image reconstruction. [22] proposed sinogram completion algorithm to address the problem in CT acquisition process. All of these tasks where our goal is to translate images from source domain to the target domain, can be considered as the instance of medical image-to-image translation problem.
With the development of deep learning algorithms in computer vision, the CNN-based generative models has been proposed to be an effective method in medical image-to-image translation. Specifically, Generative Adversarial Networks (GANs) is one of the most popular network architecture, which consists of a generator and a discriminator. The generator predicts the synthetic data from random noise, and the discriminator learn to distinguish between real training data and the synthetic data. [43] jointly train a noise reducing generator and an adversarial discriminator to denoise the low-dose cardiac CT images. With the addition of a gradient-based loss function to the pix2pix model, [29] generated CT from MRI and 7T MRI from 3T MRI. [1] cascades a series of the encoder-decoder block in the generator and combined different loss functions a new loss MedGAN framework for PET denoising and MR motion correction.
However, GAN-based models are usually task-specific, so the method might be difficult to generalize to a new image-to-image translation task without application-specific modifications. In addition, training GANs can be challenging to train due to the need to balance between the optimization of generator and discriminator (e.g find the saddle point of min-max objective). The training is therefore susceptible to non-convergence and mode collapse.
In recent years, diffusion models have outperformed other methods in generative tasks in terms of its image quality and diversity. Diffusion models iteratively add random noise to the original images until images become a gaussian noise distribution in a Markov chain, and then remove noise in the reverse process to approximate the distribution of real image. A variety of applications, including image generation, image super-resolution, image in-painting, have been accomplished with the denoising diffusion probabilistic model and score-based model [13, 39], denoted as diffusion models. [33] have developed a uniform to image-to-image translation framework Palette for above image-to-image translation tasks[33]. The cascaded diffusion model, consisting of a sequence of super-resolution diffusion models, can generate high fidelity video given a text prompt [12]. In medical imaging, DDPMs can be used to synthesize artificial histopathology images with rare cancer subtypes to mitigate data imbalance problems for medical data [27]. [25] convert CT images into corresponding MRI images with conditional DDPM and score-matching models. The forward and backward of diffusion processes are guided by T2w MRI images to guarantee the sample quality. [10] proposed to perform brain PET image denoising with MRI as prior information for improved image quality. Diffusion models have also been employed to address low-count PET image denoising for 3D whole scans with different count levels [45].
However, iterating over a large number of steps in the reverse process significantly reduces sampling speed, which is problematic for larger images. The strategy to improve sampling speed is then proposed. Denoising diffusion implicit models accelerating sampling via non-Markovian diffusion process[38]. [28] uses fewer forward passes by learning the variance of the reverse diffusion process. Another problem when applying diffusion models to medical image translation is its incapability of maintain the structural information accurately, which can be caused by the loss of structured details during the forward diffusion process and cannot be completely recovered through the learned reverse diffusion [18]. Moreover, compared with GAN, it needs no adversarial part, so the training results might be indeterministic, which is critical for medical image diagnosis. With different random starting points, the reverse process will diverge, causing inconsistencies in the translated results.
Although both GAN architecture and diffusion models have individually exhibit their capability in medical image translation tasks, the potential of combining GAN-based model and diffusion models remains largely unexplored. In this paper, we propose a novel diffusion model framework for four different medical image-to-image translation tasks, including dual energy X-ray visualization, sparse-view CT reconstruction, T1-T2 MRI conversion, and low-count PET scans denoising. The framework is consisting of two stage: the initial prediction with GAN or CNN based model, and a cascaded multi-path short cut diffusion model for further refinement. The resultant framework achieves high performance and robustness across multiple tasks.
Our contributions are summarized as follows:
  • 1.
    A new deep learning framework combining the advantages of CNN or GAN models and diffusion models is developed for a suite of medical image translation tasks. The added diffusion model refines the preliminary results of CNN or GAN models. In addition, with the prior image obtained from CNN or GAN model, it reduces the training costs for the diffusion model and increase its training stability.
  • 2.
    A novel cascaded and multi-path sampling strategy in the reverse process of the diffusion model is introduced. With different random noises added to the CNN or GAN prior, the images are transformed in multiple paths, and the averaged results are sent to the next cascade for a new round of diffusion to further improve the performance.
  • 3.
    The quantitative and qualitative results of the proposed method and other state-of-the arts method are shown for four different medical image-to-image translation tasks. The analysis of the hyperparameters for the proposed cascaded multi-path diffusion framework is shown.
2 Related Work
Generative Models. Generative models aims to generate new data instances by learning the underlying distribution of dataset and synthesizing samples that mimicking the original data distribution. Variational Autoencoders (VAE) is one of the common methods that are used for generative tasks [19]. It is a Bayesian inference-based probablistic model consisting of a encoder to learn features and representations of data in latent space and a decoder to reconstruct the representations back into its original data space. Generative Adversarial Networks (GANs) are later introduced as the mainstream generative models, based on the interplay of a generator and a discrimator. Generative Adversarial Networks (GANs) have emerged as a prominent class of generative models, characterized by the dynamic interaction between two distinct neural networks: a generator and a discriminator. The foundational concept of GANs is predicated on the adversarial relationship between these networks. The generator network is designed to produce synthetic data instances from random noise inputs, while the discriminator network functions as a binary classifier, tasked with discerning genuine data samples from the fake ones produced by the generator.
Medical Image Translation. xxx
@@ 3.1 Cascaded Multi-path Shortcut Diffusion Model
Training: Let us denote the input image as $x$ and the translation target as $y_{0}$. For the prior image generation part, we utilized a generative network, i.e. UNet [32], that aims to predict $y_{0}$ from $x$. The network can be trained in a conditional adversarial fashion [15] using both a pixel-wise L2 loss
@@
Fig. 3: Qualitative comparison of translation results and corresponding error map from different methods. Examples from DE X-ray soft-tissue generation (Left), Sparse-view CT reconstruction (Middle), and MRI T1-to-T2 synthesis are shown. The image quality metrics of each sample are indicated at the bottom left of the images.
Inference: Once the prior image generation network $f_{prior}(\cdot)$ of cGAN and the conditional diffusion network $f_{dm}(\cdot)$ are converged from training, we can use them in CMDM for image translation. The overall inference pipeline of CMDM is illustrated in Figure 2. Instead of starting the reverse process from a standard normal distribution $\mathcal{N}(y_{T}|0,I)$ at $T$, the reverse process starts at a pre-defined time point $t_{s}\in[0,T]$ with
@@
UNet $39.76\pm 2.36$ $0.984\pm 0.003$ $0.606\pm 0.071$ $41.33\pm 3.18$ $0.988\pm 0.003$ $0.571\pm 0.066$ 0.013
cGAN $39.82\pm 2.37$ $0.985\pm 0.003$ $0.603\pm 0.072$ $41.36\pm 3.17$ $0.988\pm 0.003$ $0.572\pm 0.065$ 0.013
Palette v1 $42.89\pm 2.34$ $0.987\pm 0.002$ $0.390\pm 0.047$ $43.06\pm 3.16$ $0.989\pm 0.002$ $0.373\pm 0.042$ 13.670
Palette v2 $43.11\pm 2.34$ $0.988\pm 0.002$ $0.382\pm 0.045$ $43.47\pm 3.13$ $0.990\pm 0.002$ $0.363\pm 0.043$ 273.420
I2SB $43.18\pm 2.35$ $0.988\pm 0.002$ $0.381\pm 0.045$ $43.49\pm 3.14$ $0.990\pm 0.002$ $0.367\pm 0.043$ 14.551
BBDM $43.08\pm 2.35$ $0.988\pm 0.002$ $0.382\pm 0.044$ $43.52\pm 3.13$ $0.989\pm 0.002$ $0.359\pm 0.043$ 15.121
Ours $\mathbf{44.27\pm 2.33}$ $\mathbf{0.991\pm 0.002}$ $\mathbf{0.369\pm 0.041}$ $\mathbf{44.58\pm 3.16}$ $\mathbf{0.992\pm 0.003}$ $\mathbf{0.348\pm 0.038}$ 154.663
@@
UNet $44.11\pm 1.38$ $0.977\pm 0.004$ $0.372\pm 0.047$ $46.32\pm 1.27$ $0.981\pm 0.004$ $0.315\pm 0.040$ 0.006
cGAN $44.13\pm 1.39$ $0.978\pm 0.004$ $0.370\pm 0.047$ $46.35\pm 1.28$ $0.981\pm 0.004$ $0.314\pm 0.040$ 0.006
Palette v1 $44.96\pm 1.24$ $0.980\pm 0.003$ $0.321\pm 0.041$ $46.75\pm 1.26$ $0.987\pm 0.004$ $0.310\pm 0.039$ 8.863
Palette v2 $45.56\pm 1.24$ $0.981\pm 0.003$ $0.318\pm 0.040$ $46.95\pm 1.25$ $0.988\pm 0.003$ $0.308\pm 0.038$ 177.202
I2SB $45.86\pm 1.26$ $0.982\pm 0.003$ $0.317\pm 0.039$ $46.91\pm 1.26$ $0.989\pm 0.003$ $0.309\pm 0.039$ 9.561
BBDM $45.73\pm 1.24$ $0.981\pm 0.003$ $0.318\pm 0.040$ $46.96\pm 1.26$ $0.989\pm 0.003$ $0.309\pm 0.038$ 9.987
Ours $\mathbf{46.42\pm 1.22}$ $\mathbf{0.986\pm 0.003}$ $\mathbf{0.302\pm 0.039}$ $\mathbf{47.02\pm 1.25}$ $\mathbf{0.990\pm 0.003}$ $\mathbf{0.299\pm 0.038}$ 108.821
@@
UNet $27.17\pm 1.56$ $0.885\pm 0.042$ $0.222\pm 0.051$ $27.38\pm 1.59$ $0.891\pm 0.046$ $0.216\pm 0.052$ 0.006
cGAN $27.19\pm 1.58$ $0.887\pm 0.044$ $0.220\pm 0.052$ $27.41\pm 1.58$ $0.891\pm 0.047$ $0.217\pm 0.053$ 0.006
Palette v1 $27.52\pm 1.57$ $0.890\pm 0.044$ $0.218\pm 0.051$ $27.68\pm 1.54$ $0.897\pm 0.046$ $0.210\pm 0.052$ 8.863
Palette v2 $27.68\pm 1.55$ $0.891\pm 0.043$ $0.211\pm 0.051$ $27.79\pm 1.52$ $0.899\pm 0.044$ $0.206\pm 0.051$ 177.202
I2SB $27.85\pm 1.56$ $0.892\pm 0.043$ $0.208\pm 0.051$ $27.89\pm 1.54$ $0.898\pm 0.043$ $0.208\pm 0.052$ 9.561
BBDM $27.88\pm 1.56$ $0.892\pm 0.043$ $0.207\pm 0.051$ $27.86\pm 1.53$ $0.899\pm 0.045$ $0.207\pm 0.051$ 9.987
Ours $\mathbf{27.93\pm 1.54}$ $\mathbf{0.898\pm 0.042}$ $\mathbf{0.202\pm 0.051}$ $\mathbf{27.98\pm 1.54}$ $\mathbf{0.901\pm 0.044}$$\mathbf{.201\pm.051}$ 108.821
@@ 3.2 Dataset Preparation
Fig. 4: Ablative studies on the reverse starting time (Left), the number of paths (Middle), and the number of cascades (Right). DE X-ray soft-tissue image generation and 1/6 SVCT reconstruction were utilized for these studies. Peak performances were annotated on the plots with the corresponding image quality metric, i.e. PSNR.
@@ 4 Experimental Results
Figure 3 shows qualitative comparisons between previous state-of-the-art and our methods. Examples from the DE X-ray dataset, SVCT reconstruction dataset, and MRI translation dataset are illustrated. For the DE X-ray example (left two columns), we can see all the previous translation methods can generate reasonable soft-tissue images, i.e. rib-suppression images, from the standard X-ray image. While cGAN could generate visually plausible results with a PSNR of $44.74$dB, the translated images still suffer from relatively inaccurate quantification as indicated by the error map. On the other hand, we can see the previous diffusion-based methods, e.g. Palette and BBDM, both achieved significantly better translation as compared to cGAN with PSNR improving to $46.05$dB, with much fewer pixel-wise errors indicated by the error maps. In the last row, we can find that our CMDM further improved over the previous diffusion-based methods with PSNR reaching to $47.34$dB, where further reduced pixel-wise error can be found in the cardiac and lung regions. Similarly, for the SVCT example (the middle two columns), cGAN can reasonably suppress the streak artifact in the input FBP SVCT image. However, significant residual errors can be found in the femoral head and pelvic bone regions. On the other hand, we observe that the previous diffusion-based methods can suppress these errors, with PSNR reaching close to $46$dB. Furthermore, with CMDM combining cGAN and Diffusion, we can see that the overall error of our translation results are reduced even more, and the image quality is enhanced to PSNR of $46.23$dB. Similar observations can be found for the T1-to-T2 translation example in the last two columns.
@@
The quantitative comparisons were summarized in Table 1. Similar to the observations from the visualizations, we can see the traditional CNN-based approaches generally under-performed the diffusion-based approaches. For example, the cGAN only achieved an average PSNR of $39.82$dB and MAE of $0.603$ for the soft-tissue image translation, while the single reverse path Palette, i.e. Palette v1, significantly outperformed it with PSNR of $42.89$dB and MAE of $0.390$. Running multiple reverse paths of Palette and averaging the outputs, i.e. Palette v2, led to improved performance which reached similar performances of I2SB and BBDM with PSNR of $43.11$dB and MAE of $0.382$. In the last row, our CMDM achieved an average PSNR of $44.27$dB and MAE of $0.369$ that significantly outperformed all the previous baseline methods. Comparing the soft-tissue image translation task to the bone image translation task, the CMDM had slightly higher performance on the latter task since the bone image without complex soft-tissue texture can be relatively easier to generate as compared to the soft-tissue image. For the inference speed in DE X-ray applications, I2SB and BBDM with a single reverse path took an average of $14.55$ and $15.12$ seconds, respectively. CMDM with the best performance took an average of $154.66$ seconds per inference since multiple shortcut reverse paths are needed. Similar to the quantitative results for the DE X-ray, we found our CMDM consistently outperformed previous CNN and diffusion-based baseline methods for both the SVCT reconstruction applications and the MRI inter-modality translation applications.
@@
Fig. 5: Examples of CMDM’s uncertainty estimation for DE X-ray soft-tissue image generation (Left) and MRI T1-to-T2 synthesis (Right). The relationship plots between the absolute error (bottom left) and the uncertainty (bottom right) were shown as well. Positive correlations with $R>0.75$ were found for both cases.
Table 2: Quantitative comparison of CMDM with different prior strategies. Analysis with DE X-ray soft-tissue generation task, 1/6 SVCT reconstruction task, and T1-to-T2 MRI synthesis task are reported.
MAE DE X-Ray SVCT MRI
w/o prior $0.379\pm 0.043$ $0.316\pm 0.042$ $0.210\pm 0.051$
UNet prior $0.370\pm 0.041$ $0.306\pm 0.041$ $0.203\pm 0.051$
UFNet prior $0.366\pm 0.041$ $0.303\pm 0.040$ $0.201\pm 0.050$
cGAN prior $0.369\pm 0.041$ $0.302\pm 0.039$ $0.201\pm 0.051$
@@
We conducted ablative studies for the hyper-parameters in CMDM, including the reverse starting time, the number of shortcut paths, and the number of cascades. The results for the DE X-ray and SVCT are summarized in Figure 4. First, for the reverse starting time, we can see that setting $t_{s}$ to around $200$ yields the best performance, and the performance starts to degrade if we further increase it. It is worth noticing that using $t_{s}=200$ here not only yields the best performance but allows us to reduce the inference time by about 5 times as compared to the previous diffusion methods that start at $t=1000$ or beyond. Second, for the number of shortcut paths, we can see that the performance increases as we use an increasing number of paths. The performance started to converge when $20$ paths were used. Because the inference time increased linearly as we increased the number of paths, we chose the converging point $N_{p}=20$ in our method. Thirdly, for the number of cascades, we found that the performance gradually boosted as the number of cascades increased. However, peak performance was reached when $N_{c}=3$, and the inference started to overfit, leading to degraded translation performance. Lastly, we investigated the impact on CMDM when different prior image generations were used, including priors from UNet [32], Under-to-fully-complete Network (UFNet [51]), and cGAN [15]. As we can see from Table 2, using CMDM with prior always outperforms CMDM without prior. Among all the prior generated, CMDM with priors generated from cGAN and UFNet yields the best performance. Moreover, we also studied CMDM with or without the conditional input for the diffusion part. As we can see from Table 3, CMDM without conditional input can still generate a reasonable translation guided by the prior image. However, CMDM with condition input with more translation guidance still yields the best performance.
Table 3: Quantitative comparison of CMDM with or without images to be translated as conditional inputs in the diffusion part. Analysis with DE X-ray soft-tissue generation task, 1/6 SVCT reconstruction task, and T1-to-T2 MRI synthesis task are reported.
MAE DE X-Ray SVCT MRI
w/o condition $0.517\pm 0.059$ $0.339\pm 0.043$ $0.219\pm 0.053$
w condition $0.369\pm 0.041$ $0.302\pm 0.039$ $0.202\pm 0.051$